# Topological advantage for adsorbate chemisorption on conjugated chains

### v2

@@ III.2 Chemisorption on solitons and trivial defects
In an ideal SSH chain, midgap ($E\!\approx\!0$) modes occur only at the edges of the topological phase $r>1$. At finite temperature, bulk topological defects known as solitons (domain walls) [su1979solitons, takayama1980continuum, heeger1988solitons] arise and bind additional exponentially localized midgap states in the interior. These localized states may also hybridize efficiently with nearby molecular orbitals, enabling charge donation to an adsorbate. In contrast, a metallic SSH chain ($r=1$) lacks localized midgap modes, but has greater density of states of near-resonant levels with the adsorbate, and in the presence of trivial defects (e.g., impurities) may also have local resonances that could hybridize effectively with adsorbates. We therefore compare multi-adsorbate chemisorption on the following substrates: (i) a topological chain with multiple solitons (soliton ensemble), (ii) an ordered metallic chain, and (iii) a disordered metallic chain with trivial defects.
@@
We report the average LUMO occupancy $\langle n_{0}\rangle$ [Eq. (8)] for $N_{\text{ad}}$ adsorbates on a chain of $N$ unit cells at 0 K and half filling. Domain walls have width $\xi\in\{5,7,10\}$ (corresponding to the localization length of their midgap states, see Eq. (6)). In the soliton ensemble simulations reported in Fig. 5a), each adsorbate sits at a soliton center and $N_{\text{ad}}$ equals the number of domain walls $N_{DW}$ (including the two edges). In the metallic phase with trivial defects, we set $r=1$ and introduce a dilute ensemble of defects by drawing small near-zero on-site potentials from a Gaussian distribution, applied to a random subset of sites to mimic weak non-topological disorder. Adsorbates are then placed at defect locations drawn from a uniform distribution over bulk sites. Finally, in the clean metal scenario, $r=1$ and the adsorbates are placed at uniformly random bulk sites. Unless otherwise specified, these simulations employed $N=3000$, $v=10$, $T_{m}=0.1$, and averages over 25 realizations with fixed electron number.
@@
Figure 5: Resonant case $\varepsilon_{0}=0$. (a) $\braket{n_{0}}$ vs. number of adsorbates (defects) for soliton ensembles at three dimerizations ($r=1.5,1.1,1.0001$), compared to (i) metal ($r=1$ with the same number of on-site defects and (ii) clean metal ($r=1$, no defects). Shaded regions represent fluctuations over realizations of the disordered systems, and we have set the number of adsorbates to be equal the number of domain walls $N_{ad}=N_{DW}$. (b) $\braket{n_{0}}$ vs. $N_{\text{ad}}$ for a fixed set of 40 domain walls (topological) or none (clean metal), and adsorbates placed at random sites. Parameters: $\xi=7$, $N=3000$, $v=10$, $T_{m}=0.1$, and 25 realizations.
Figure 6: Off-resonant case $\varepsilon_{0}=0.1$: $\braket{n_{0}}$ vs. $N_{\text{ad}}$ and defect content. Parameters: $N=3000$, $v=10$, $T_{m}=0.1$.
@@
Across all conditions, chains with domain walls yield systematically larger $\braket{n_{0}}$ than metallic chains with trivial defects (Fig. 5a). The enhancement tracks soliton localization: larger $r$ (stronger dimerization) narrows midgap wavefunctions and increases adsorbate–chain overlap, consistent with SSH theory [su1979solitons]. Near criticality ($r=1.0001$) midgap states broaden and charge transfer is reduced. Fluctuations reflect variability from soliton–antisoliton spacing and boundary proximity.
@@
The near plateau $\braket{n_{0}}\approx 0.25$ at $r=1.1$ (Fig. 5a) can be explained as follows. At half filling and with particle–hole symmetry, each isolated domain wall binds one midgap state at zero energy which is half filled on average. Given that any adsorbate positioned in the neighborhood of a particular soliton can reach at most $n_{0}\approx 0.5$ when $\varepsilon_{0}\geq 0$ (Fig. 3), averaging over all adsorbates gives a low-coverage plateau of $\langle n_{0}\rangle\approx(0.5)\times(0.5)=0.25$. As the number of domain walls increases, domain-wall interactions lead to the observed decay in $\langle n_{0}\rangle$. Metals ($r=1$) with trivial defects behave similarly to clean metals, i.e., both give much lower $\braket{n_{0}}$ relative to the system with topological defects, thus indicating that non-topological near-zero resonances do not generically replicate midgap-assisted donation and instead hybridize significantly with metal extended states rather than to the adsorbate orbital.
@@
Figure 5b fixes forty solitons but randomizes adsorbate positions, hence, the adsorbates may or may not land on a domain wall. At low coverage the same plateau $\langle n_{0}\rangle\!\approx\!0.25$ emerges, but with large variance: realizations in which adsorbates overlap midgap states yield $n_{0}\!\sim\!0.5$, whereas those on bulk sites yield $n_{0}\!\sim\!0$. This two-population mixture (soliton-bound vs. bulk-bound) produces both the $0.25$ mean and the broad fluctuations (shaded blue region). In contrast, clean metallic chains give substantially lower mean LUMO occupancy at all coverages, underscoring the chemisorption advantage conferred by topological midgap states.
@@
Off resonance ($\varepsilon_{0}>0$; Fig. 6) the topological advantage persists but is reduced. The dependence on $\xi$ is weak-to-moderate: narrower domain walls (smaller $\xi$) produce more localized midgap states and slightly larger $\langle n_{0}\rangle$ due to improved wavefunction overlap. In contrast, the trivial-defect metal behaves similarly to the ordered metal, indicating that non-topological near-zero local resonances do not generically replicate the midgap-assisted donation of domain walls.
III.3 Topological phase transition signature on adsorbate electronic friction
@@
Using the adiabatic orbitals from diagonalizing Eq. (2) with $\varepsilon_{0}(R)=\epsilon_{d}+R\,g\sqrt{m\omega/\hbar}$, we construct the local molecular density of states $P(\epsilon,R)$ and evaluate the electronic friction $\gamma(R)$ via Eq. (II.2) for the metallic ($r=1$), topological ($r>1$), and trivial ($r<1$) phases at temperature $1/\beta=0.015$. As explained in Sec. II, $P(\epsilon,R)$ is obtained by Gaussian broadening of each adiabatic level $\epsilon_{\lambda}(R)$ with bandwidth $\sigma=0.019$. While $\sigma$ is not uniquely defined [maurer2016ab], this choice yields smooth non-oscillatory $P(\epsilon,R)$ and $\gamma(R)$ sufficient for qualitative trends across phases (see Appendix A for a bandwidth sensitivity analysis). For a dilute impurity concentration and fixed electron number in the polyacetylene chain, we treat the system effectively as intrinsic [ashcroft2022solid]. We modeled the interacting adsorbate as a hole-like impurity and set the Fermi level $\mu$ in the HOMO–LUMO midgap at half filling (spinless electrons, for $N$ unit cells). Because hybridization shifts the spectrum, $\mu$ is tracked self-consistently as a function of $R$; see Fig. 8b.
@@
Figure 7: Electronic friction $\gamma(R)$ [Eq. (II.2)] across phases. (a) Topological edge ($x_{E}$; purple, $r=1.1$), metallic bulk ($x_{B}$; magenta, $r=1$), and trivial bulk ($x_{B}$; teal, $r=0.9$) at $T_{1,A}=T_{N/2,A}=0.1$. (b) Same as (a) with weaker coupling $T_{1,A}=T_{N/2,A}=0.01$. (c,d) Near criticality: topological edge ($x_{E}$; green, $r=1.01$) and trivial bulk ($x_{B}$; orange, $r=0.99$) vs. $\varepsilon_{0}(R)$ at $T_{1,A}=0.1$ (c) and $T_{1,A}=0.01$ (d). Parameters: $v=10$, $\sigma=0.019$, $\varepsilon_{d}=0.15$, $\kappa=0.02$, $\beta=1/0.015$, $N=800$, with $T_{N/2-1,B}=T_{N/2,B}=T_{N/2,A}/3$.
@@
Figure 7a shows $\gamma(R)$ vs. $\varepsilon_{0}(R)$ for representative placements and dimerizations. The metallic phase exhibits the largest friction, consistent with a high LDOS at $\mu$ and abundant low-energy electron–hole excitations. In the topological phase at the edge, the localized boundary mode enhances friction relative to a gapped bulk but remains below the metallic value. The reduction arises from the hybridization-induced splitting between the edge mode and the LUMO: a two-level avoided crossing produces a PDOS doublet $\langle\varepsilon_{0}|P(\epsilon,R)|\varepsilon_{0}\rangle$ separated by $\Delta E\sim 2\mathcal{O}(T_{1,A})$, which depresses the LDOS at $\mu$ and suppresses e–h generation (Fig. 8a). At smaller coupling (Fig. 7b), $\Delta E$ decreases, restoring LDOS at $\mu$ and increasing $\gamma(R)$.
At smaller adsorbate–edge coupling $T_{1,A}=0.01$ [Fig. 7b], the splitting $\Delta E$ diminishes, partially restoring LDOS at $E_{F}$ and increasing $\gamma(R)$ relative to the stronger-coupling case. Appendix A details the evolution of the PDOS doublet with $T_{1,A}$ and its correlation with the friction peak height and position. As shown there, the splitting results in a low local density of states at the Fermi energy, suppressing the electronic friction by limiting the availability of low-lying electronic transitions for electron-hole pair formation near the Fermi level.
@@
For a sizeable SSH gap |wv|=1|w-v|=1 (e.g., $v=10$, $w=9$), thermal carriers are exponentially suppressed and the adsorbate level lies far from the valence/conduction continua [Fig. 8b]. With $\mu$ pinned near midgap, $\gamma(R)$ is negligible in the trivial phase, consistent with the absence of low-energy e–h channels. Likewise, in the topological phase but in the bulk (away from edges), $\gamma(R)$ is similarly suppressed; we therefore omit those curves from Fig. 7 for clarity. In such gapped settings, vibrational energy relaxation is dominated by phonon-mediated channels in the insulating substrate [sjakste2018energy, giustino2017electron].
@@ IV Conclusions
We find a robust advantage for electron donation in the topological phase relative to both the metallic and trivial phases. This is explained by the fact that exponentially localized midgap states at edges (and at solitons) have large wavefunction amplitude at the adsorption site, which enhances adsorbate–substrate hybridization at resonance. By contrast, although the metallic phase has a larger total density of states near the Fermi level, its extended bulk states are spatially dilute at any one site, weakening local hybridization and reducing charge transfer. Importantly, merely generating localized states is necessary but not sufficient: the donation depends sensitively on resonance alignment, spatial overlap, and occupancy. Topological midgap modes satisfy these criteria in a robust, symmetry-protected manner, whereas trivial near-zero resonances in a metal generally do not. Consistent with this picture, solitons in the topological phase substantially outperform trivial defects in a metal. Near criticality (weak dimerization regime $r\approx 1$) the midgap states broaden and the advantage correspondingly diminishes.
@@
Appendix A Electronic friction bandwidth
In our computations of electronic friction the discrete spectrum entering the local (projected) density of states is broadened by replacing Dirac distributions with Gaussians of width $\sigma$ [Eq. (14)], yielding a smooth LDOS $P(\epsilon,R)$. This numerical broadening is chosen to (i) resolve physical features while (ii) avoiding spurious oscillations when evaluating the electronic–friction integrand, which involves the overlap of $P(\epsilon,R)$ with the thermal factor $\partial_{\epsilon}f_{T}(\epsilon)$ [cf. Eq. (II.2)]. Throughout the main text we use $\sigma=0.019$, set by the single–particle level spacing of the bare SSH chain near $E_{F}$ in the metallic case for $N=800$ unit cells; see Fig. 9 for a sensitivity analysis.
Figure 9 illustrates how $\sigma$ affects $\gamma(R)$. In the metal ($r=1$, bulk adsorbate), smaller $\sigma$ sharpens the LDOS at $E_{F}$, increasing the friction peak. Larger $\sigma$ smears spectral weight away from $E_{F}$ and reduces $\gamma$. In the topological phase at the edge ($r=1.1$), $\gamma(R)$ shows much weaker dependence on $\sigma$ because the HOMO–LUMO gap and edge–adsorbate hybridization minimize the overlap of $P(\epsilon,R)$ with $\partial_{\epsilon}f_{T}$ except near the avoided crossing.
Figure 9: Dependence of electronic friction $\gamma(R)$ [Eq. (II.2)] on Gaussian broadening $\sigma$ in the topological phase at the edge ($r=1.1$, $x_{E}$) and in the metallic phase in the bulk ($r=1$, $x_{B}$). Two values are shown: $\sigma=0.019$ and $\sigma=0.038$. Other parameters are $v=10$, $g=0.02$, $\beta=1/0.015$, $N=800$, and adsorbate couplings $T_{1,A}=T_{N/2,A}=0.1$ with $T_{N/2-1,B}=T_{N/2,B}=T_{N/2,A}/3$.
In the topological edge case, the leading effect of the molecule–boundary coupling $T_{1,A}$ is an avoided crossing between the edge mode and the adsorbate level, producing a PDOS doublet in $\langle\varepsilon_{0}|P(\epsilon,R)|\varepsilon_{0}\rangle$ with splitting $\Delta E\sim 2|T_{1,A}\alpha_{1}^{L}|$. When $\Delta E$ is large (e.g., $T_{1,A}=0.1$), the Fermi level $E_{F}$ falls between the two PDOS peaks, reducing the overlap with $\partial_{\epsilon}f_{T}$ and thus suppressing $\gamma(R)$. As $T_{1,A}$ decreases, the peaks approach and merge near $E_{F}$, increasing the overlap and $\gamma(R)$. Importantly, in a finite topological chain one edge remains weakly or non-interacting; with fixed particle number the resulting spectral rearrangements keep $E_{F}$ near zero as $T_{1,A}\!\to\!0$.
Figure 10 compares the PDOS on the adsorbate at $\varepsilon_{0}=0$ for the topological edge and metallic bulk placements, together with $\partial_{\epsilon}f_{T}$ on a secondary axis to visualize the overlap that controls friction. In the metal a single PDOS peak aligns with $E_{F}$, maximizing overlap. At the topological edge the split peaks flank $E_{F}$, reducing overlap—unless the coupling is weak enough that the doublet collapses.
Figure 10: Projected DOS on the adsorbate LUMO, $\langle\varepsilon_{0}|P(\epsilon,R)|\varepsilon_{0}\rangle$, at $\varepsilon_{0}=0$ for the topological edge (solid) and metallic bulk (dotted), plotted together with $\partial_{\epsilon}f_{T}(\epsilon)$ (right axis) to show the overlap relevant to friction. The employed parameters were $v=10$, $T_{1,A}=T_{N/2,A}=0.1$, $\beta=1/0.015$, $N=800$, $\sigma=0.019$, and $T_{N/2-1,B}=T_{N/2,B}=T_{N/2,A}/3$.
Across the tested $\sigma$ values, our qualitative conclusions are unchanged: (i) metallic friction exceeds topological-edge friction due to larger LDOS at $E_{F}$; (ii) trivial (and bulk-topological) gapped regions yield negligible friction; and (iii) weaker molecule–edge coupling reduces hybridization splitting and increases friction by restoring PDOS overlap at $E_{F}$.

### v3

@@ Abstract
Chemisorption on solitons and trivial defects In an ideal SSH chain, midgap ($E\!\approx\!0$) modes occur only at the edges of the topological phase $r>1$. At finite temperature, bulk topological defects known as solitons (domain walls) su1979solitons, takayama1980continuum, heeger1988solitons arise and bind additional exponentially localized midgap states in the interior. These localized states may also hybridize efficiently with nearby molecular orbitals, enabling charge donation to an adsorbate. In contrast, a metallic SSH chain ($r=1$) lacks localized midgap modes, but has greater density of states of near-resonant levels with the adsorbate, and in the presence of trivial defects (e.g., impurities) may also have local resonances that could hybridize effectively with adsorbates. We therefore compare multi-adsorbate chemisorption on the following substrates: (i) a topological chain with multiple solitons (soliton ensemble), (ii) an ordered metallic chain, and (iii) a disordered metallic chain with trivial defects.
@@
We report the average LUMO occupancy $\langle n_{0}\rangle$ [Eq. (8)] for $N_{\text{ad}}$ adsorbates on a chain of $N$ unit cells at 0 K and half filling. Domain walls have width $\xi\in\{5,7,10\}$ (corresponding to the localization length of their midgap states, see Eq. (6)). In the soliton ensemble simulations reported in Fig. 5a), each adsorbate sits at a soliton center and $N_{\text{ad}}$ equals the number of domain walls $N_{DW}$ (including the two edges). In the metallic phase with trivial defects, we set $r=1$ and introduce a dilute ensemble of defects by drawing small near-zero on-site potentials from a Gaussian distribution, applied to a random subset of sites to mimic weak non-topological disorder. Adsorbates are then placed at defect locations drawn from a uniform distribution over bulk sites. Finally, in the zero disroder metal scenario, $r=1$ and the adsorbates are placed at uniformly random bulk sites. Unless otherwise specified, these simulations employed $N=3000$, $v=10$, $T_{m}=0.1$, and averages over 25 realizations with fixed electron number.
@@
Figure 5: Resonant case $\varepsilon_{0}=0$. (a) $\expectationvalue{n_{0}}$ vs. number of adsorbates (defects) for soliton ensembles at three dimerizations ($r=1.5,1.1,1.0001$), compared to (i) metal ($r=1$ with the same number of on-site defects and (ii) disorder-free metal ($r=1$, no defects). Shaded regions represent fluctuations over realizations of the disordered systems, and we have set the number of adsorbates to be equal the number of domain walls $N_{ad}=N_{DW}$. (b) $\expectationvalue{n_{0}}$ vs. $N_{\text{ad}}$ for a fixed set of 40 domain walls (topological) or none (zero-disorder metal), and adsorbates placed at random sites. Parameters: $\xi=7$, $N=3000$, $v=10$, $T_{m}=0.1$, and 25 realizations.
[…]
@@
Across all conditions, chains with domain walls yield systematically larger $\expectationvalue{n_{0}}$ than metallic chains with trivial defects (Fig. 5a). The enhancement tracks soliton localization: larger $r$ (stronger dimerization) narrows midgap wavefunctions and increases adsorbate–chain overlap su1979solitons. Near criticality ($r=1.0001$) midgap states broaden and charge transfer is reduced. Fluctuations reflect variability from soliton–antisoliton spacing and boundary proximity.
@@
The near plateau $\expectationvalue{n_{0}}\approx 0.25$ at $r=1.1$ (Fig. 5a) can be explained as follows. At half filling and with particle–hole symmetry, each isolated domain wall binds one midgap state at zero energy which is half filled on average. Given that any adsorbate positioned in the neighborhood of a particular soliton can reach at most $n_{0}\approx 0.5$ when $\varepsilon_{0}\geq 0$ (Fig. 3), averaging over all adsorbates gives a low-coverage plateau of $\expectationvalue{n_{0}}\approx(0.5)\times(0.5)=0.25$. As the number of domain walls increases, domain-wall interactions lead to the observed decay in $\expectationvalue{n_{0}}$. Metals ($r=1$) with trivial defects behave similarly to clean metals, i.e., both give much lower $\expectationvalue{n_{0}}$ relative to the system with topological defects, thus indicating that non-topological near-zero resonances do not generically replicate midgap-assisted donation and instead hybridize significantly with metal extended states rather than to the adsorbate orbital.
@@
Figure 5b fixes forty solitons but randomizes adsorbate positions, hence, the adsorbates may or may not land on a domain wall. At low coverage the same plateau $\expectationvalue{n_{0}}\!\approx\!0.25$ emerges, but with large variance: realizations in which adsorbates overlap midgap states yield $n_{0}\!\sim\!0.5$, whereas those on bulk sites yield $n_{0}\!\sim\!0$. This two-population mixture (soliton-bound vs. bulk-bound) produces both the $0.25$ mean and the broad fluctuations (shaded blue region). In contrast, ordered metallic chains give substantially lower mean LUMO occupancy at all coverages, underscoring the chemisorption advantage conferred by topological midgap states.
@@
Off resonance ($\varepsilon_{0}>0$; Fig. 6) the topological advantage persists but is reduced. The dependence on $\xi$ is weak-to-moderate: narrower domain walls (smaller $\xi$) produce more localized midgap states and slightly larger $\expectationvalue{n_{0}}$ due to enhanced wavefunction overlap. In contrast, the trivial-defect metal behaves similarly to the ordered metal, indicating that non-topological near-zero local resonances do not generically replicate the midgap-assisted donation of domain walls.
@@
Topological phase transition signature on adsorbate electronic friction. Using the adiabatic orbitals from diagonalizing Eq. (2) with $\varepsilon_{0}(R)=\epsilon_{d}+R\,g\sqrt{m\omega/\hbar}$, we construct the local molecular density of states $P(\epsilon,R)$ and evaluate the electronic friction $\gamma(R)$ via Eq. (Topological advantage for adsorbate chemisorption on conjugated chains) for the metallic ($r=1$), topological ($r>1$), and trivial ($r<1$) phases at temperature $1/\beta=0.015$. As explained above, $P(\epsilon,R)$ is obtained by Gaussian broadening of each adiabatic level $\epsilon_{\lambda}(R)$ with width $\sigma=0.0225$. While $\sigma$ is not uniquely defined maurer2016ab, this choice yields smooth non-oscillatory $P(\epsilon,R)$ and $\gamma(R)$ sufficient for qualitative trends across phases (see Supporting Information for a sensitivity analysis). For a dilute impurity concentration and fixed electron number in the polyacetylene chain, we treat the system effectively as intrinsic ashcroft2022solid. We modeled the interacting adsorbate as a hole-like impurity and set the Fermi level $\mu$ in the HOMO–LUMO midgap at half filling (spinless electrons, for $N$ unit cells). Because hybridization shifts the spectrum, $\mu$ is tracked self-consistently as a function of $R$; see Fig. 8b.
@@
Figure 7: Electronic friction $\gamma(R)$ [Eq. (Topological advantage for adsorbate chemisorption on conjugated chains)] across phases. (a) Topological edge ($x_{E}$; purple, $r=1.1$), metallic bulk ($x_{B}$; magenta, $r=1$), and trivial bulk ($x_{B}$; teal, $r=0.9$) at $T_{1,A}=T_{N/2,A}=0.1$. (b) Same as (a) with weaker coupling $T_{1,A}=T_{N/2,A}=0.01$. (c,d) Near criticality: topological edge ($x_{E}$; green, $r=1.01$) and trivial bulk ($x_{B}$; orange, $r=0.99$) vs. $\varepsilon_{0}(R)$ at $T_{1,A}=0.1$ (c) and $T_{1,A}=0.01$ (d). Parameters: $v=10$, $\sigma=0.0225$, $\varepsilon_{d}=0.15$, $g=0.02$, $\beta=1/0.015$, $N=800$, with $T_{N/2-1,B}=T_{N/2,B}=T_{N/2,A}/3$.
@@
Figure 7a shows $\gamma(R)$ vs. $\varepsilon_{0}(R)$ for representative placements and dimerization $r=v/w$. The metallic phase exhibits the largest friction, consistent with a high LDOS at $\mu$ and abundant low-energy electron–hole excitations. In the topological phase at the edge, the localized boundary mode enhances friction relative to a gapped bulk but remains below the metallic value. The reduction arises from the hybridization-induced splitting between the edge mode and the LUMO: a two-level avoided crossing produces a PDOS doublet $\langle\varepsilon_{0}|P(\epsilon,R)|\varepsilon_{0}\rangle$ separated by $\Delta E\sim 2\mathcal{O}(T_{1,A})$, which depresses the LDOS at $\mu$ and suppresses electron-hole pair generation (Fig. 8a) near the Fermi level.
At smaller adsorbate–edge coupling $T_{1,A}=0.01$ [Fig. 7b], the splitting $\Delta E$ diminishes, partially restoring LDOS at $E_{F}$ and increasing $\gamma(R)$ relative to the stronger-coupling case. Note this enhancement occurs as one reduces a strong coupling. For sufficiently weak coupling, $\gamma$ ultimately decreases with $T_{1,A}$ as the adsorbate decouples (see Supporting Information for further discussion).
@@
For a sizeable SSH gap |wv|=1|w-v|=1 (e.g., $v=10$, $w=9$), thermal carriers are exponentially suppressed and the adsorbate level lies far from the valence-conduction continua [Fig. 8b]. With $\mu$ pinned near midgap, $\gamma(R)$ is negligible in the trivial phase, consistent with the absence of low-energy electron-hole channels. Likewise, in the topological phase but in the bulk (away from edges), $\gamma(R)$ is similarly suppressed; we therefore omit those curves from Fig. 7 for clarity. In such gapped settings, vibrational energy relaxation is likely dominated by phonon-mediated channels in the insulating substrate sjakste2018energy, giustino2017electron.
@@
We find a robust advantage for electron donation in the topological phase relative to both the metallic and trivial phases. This is explained by the fact that exponentially localized midgap states at edges (and at solitons) have large wavefunction amplitude at the adsorption site, which enhances adsorbate–substrate hybridization at resonance. By contrast, although the metallic phase has a larger total density of states near the Fermi level, its extended bulk states are spatially dilute at any one site, weakening local hybridization and reducing charge transfer. Importantly, merely generating localized states is necessary but not sufficient: the donation depends also on resonance alignment, spatial overlap, and occupancy. Topological midgap modes satisfy these criteria in a robust, symmetry-protected manner, whereas trivial near-zero resonances in a metal generally do not. Consistent with this picture, solitons in the topological phase substantially outperform trivial defects in a metal. Near criticality (weak dimerization regime $r\approx 1$) the midgap states broaden and the advantage correspondingly diminishes.
@@
Finally, our choice of a spinless, weakly correlated SSH model ensures our modeling captures symmetry-protected midgap orbitals at edges and domain walls and their hybridization with an adsorbate. In more complex topological materials of interest for catalysis, a detailed treatment of spin degrees of freedom, strong spin–orbit coupling, and strong electron–electron correlations may reshape the low-energy spectrum, change the relevant symmetry class, and open additional relaxation channels. Extending the present framework to spinful tight-binding models (including SOC) and to beyond-mean-field correlated descriptions (e.g., SSH–Hubbard-type models) is an important direction for quantifying the robustness of the trends identified here and for exploring regimes where electronic correlations qualitatively alter charge transfer and dissipation at molecule–surface interfaces.
Acknowledgments R.F.R. acknowledges generous start-up funds from Emory University.
{suppinfo}
We provide evidence of lack of size dependence of our results by reporting LUMO electronic occupation numbers and electronic friction versus SSH chain length $N$ and discuss the Gaussian broadening $\sigma$ used in the friction calculation, including its effect on the adsorbate-projected DOS in the trivial and topological phases.